OpenStudy (anonymous):

what is equivalent to 2(x-5)+4x-18=4x

OpenStudy (anonymous):

2x-10+4x-18-4x=0 2x-10+4x-4x-18=0 2x-10-18=0 2x-28=0 2x=28 x=28/2 x=14

OpenStudy (anonymous):

thats wrong dummy

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i said equivalent

OpenStudy (anonymous):

sorry i thought you were solving for x, i didnt read the question

OpenStudy (anonymous):

im sorrry tooo

OpenStudy (anonymous):

6x-28=4x or 2(3x-14)=4x

OpenStudy (anonymous):

i don't know if that's the answer you are looking for